# Helios: A Foundational Language Model for Smart Energy Knowledge Reasoning and Application

### v1

@@ Abstract
In the global drive toward carbon neutrality, deeply coordinated smart energy systems underpin industrial transformation, yet their interdisciplinary, fragmented, and fast-evolving expertise prevents general-purpose large language models (LLMs), lacking domain knowledge and physical-constraint awareness, from delivering precise engineering-aligned inference and generation. To address these challenges, we introduce Helios, the first large language model tailored to the smart energy domain, together with a comprehensive suite of resources to advance LLMs research in this field. Specifically, we develop EnerSys, a multi-agent collaborative framework for end-to-end dataset construction, through which we produce: (1) the first smart energy knowledge base, EnerBase, to enrich the model’s foundational expertise; (2) the first instruction tuning dataset, EnerInstruct, to strengthen performance on domain-specific downstream tasks; and (3) the first Reinforcement Learning from Human Feedback (RLHF) dataset, EnerReinforce, to align the model with human preferences and industry standards. Leveraging these resources, Helios undergoes large-scale pretraining, instruction tuning, and RLHF. We also release EnerBench, the first benchmark for evaluating LLMs in smart energy scenarios, and demonstrate that our approach significantly enhances domain knowledge mastery, task execution accuracy, and alignment with human preferences. All training data and model checkpoints are publicly available at https://github.com/fine68/Helios.
@@
Haoyu Jiang1,†thanks: These authors contributed equally to this work.  Fanjie Zeng2,†   Boan Qu2,†   Xiaojie Lin1,∗thanks: Corresponding author.  Wei Zhong1 1 College of Energy Engineering, Zhejiang University  2 Polytechnic Institute, Zhejiang University {haoyu.jiang, zengfanjie, boan.qu, xiaojie.lin, wzhong}@zju.edu.cn https://helios-llm.github.io/
@@ 1 Introduction
Driven by the global pursuit of carbon neutrality, smart energy systems must enhance overall efficiency through the intelligent coordination of renewable energy integration, energy storage dispatch, and demand‑side response Lund et al. (2017); Dincer and Acar (2017). Smart energy is highly interdisciplinary, encompassing power engineering, information science, economics, and other fields, and its knowledge base is fragmented and rapidly evolving Ceglia et al. (2020); Thellufsen et al. (2020). Building on recent advances in general large language models (LLMs) in semantic understanding, logical reasoning, and multitask generalization, a growing body of research has used fine‑tuning and prompt engineering to adapt LLMs to task‑specific applications in smart energy, such as load forecasting Jin et al. (2023); Liao et al. (2025); Hu et al. (2025), building energy consumption modeling Wang et al. (2025b); Jiang et al. (2024), and HVAC fault diagnosis Zhang et al. (2025), thereby supporting case modeling and intelligent decision making.
@@
However, general LLMs often deliver reasoning that is semantically plausible yet physically invalid Friel and Sanyal (2023). This limitation arises chiefly because their pre‑training corpora lack reliable knowledge from the smart energy domain, leaving the models without essential domain context and physical constraints Deng et al. (2024). Current approaches Hu et al. (2025); Zhang et al. (2025) mainly invoke the prior knowledge already embedded in LLMs and do not explicitly enrich them with smart energy expertise. To alleviate these challenges, we introduce first-ever open‑sourced foundational LLM for the smart‑energy domain, referred to as Helios (Originating from the ancient Greek sun‑god, signifies the illumination of the pathway toward sustainable development through the radiance of smart energy, thereby advancing the harmonious co‑existence of humanity and the natural environment). Helios is capable of effectively tackling a broad spectrum of smart‑energy tasks. Furthermore, we present EnerSys, an end-to-end multiagent collaborative framework for dataset construction that integrates automated data generation, screening, and refinement, thereby furnishing Helios with an extensive and high‑quality data foundation.
@@
EnerSys covers three dataset‑construction phases (as shown in Fig. 1): In the construction of the pre-training dataset, the Parsing‑Agent and Deduplication Agent extract structured knowledge from the Smart Energy Corpus (scientific papers, domain‑modeling code, IEA datasets, etc.) and eliminate redundancy, building a comprehensive, balanced smart energy domain knowledge base, EnerBase; In instruction‑tuning dataset construction, on expert‑crafted seed data, we deploy Expert‑Agents for each of 14 smart‑energy sub‑domains, letting them generate instruction–response pairs from the seeds and a high‑quality corpus; the Check‑Agent then scores samples on accuracy, completeness, relevance, and usability, and the Refine‑Agent automatically fixes those below par. This pipeline yielded the EnerInstruct; In the RLHF dataset construction, agents like Write‑like‑Human craft multi‑level candidate answers to given questions, thereby creating the EnerReinforce to supply the reward model with differentiated contrastive samples. Using these datasets, we complete Helios pre‑training (adding domain basics), supervised fine‑tuning (boosting downstream skills), and RLHF reinforcement (aligning with human preferences). Concurrently, adhering to a dual‑track paradigm of “public item‑bank retrieval + expert‑targeted design,” we build EnerBench, containing 625 subjective and 976 objective questions, to systematically assess LLMs performance in smart‑energy scenarios. Experiments show Helios surpasses general‑purpose LLMs on both tasks, with output style tightly matching professional context. Our contributions can be summarized as follows:
@@ 2 Related Work
Foundation Language Models. LLMs trained on vast amounts of diverse and heterogeneous data, have accumulated extensive domain knowledge and contextual modeling capabilities. They have demonstrated human-level performance in many tasks. LLMs can be categorized into two types: 1) Closed-source models (such as OpenAI o1 Jaech et al. (2024) and Claude): These models provide inference interfaces via APIs, making them suitable for industrial-grade deployment without the need for building custom computational resources. However, they cannot be customized or extended according to specific needs; 2) Open-source models (such as DeepSeek Guo et al. (2025); Liu et al. (2024), LLaMA and Qwen Bai et al. (2023, 2025)): These models offer complete training weights, allowing for customized applications based on downstream task requirements. This has led to the development of instruction-tuned models like Alpaca Taori et al. (2023a), Vicuna Chiang et al. (2023), and Dolly Conover et al. (2023a). In this process, the quality of datasets becomes a critical factor affecting training outcomes.
@@
Domain Language Models. LLMs excel in general reasoning Jiang et al. (2025); Nam et al. (2024), their performance in specialized applications is hampered by a lack of domain expertise. This limitation has led researchers to adapt foundation models for vertical domains such as medicine Tian et al. (2024); Lin et al. (2025), chemistry Zhang et al. (2024); Zheng et al. (2025), ocean science Bi et al. (2024), and geography Deng et al. (2024). However, most domains are still in the early stages of exploration. In the energy sector, existing research primarily leverages general models’ prior knowledge through prompt engineering or fine-tuning for load forecasting Jin et al. (2023); Hu et al. (2025); Wu and Ling (2024); Wang et al. (2025a), building energy modeling Wang et al. (2025b); Jiang et al. (2024), and HVAC fault diagnosis Zhang et al. (2025). These approaches focus on application (applying LLMs’ prior knowledge to downstream tasks) rather than accumulation (enriching models with energy domain knowledge through pretraining). Furthermore, the energy domain faces a scarcity of high-quality training data due to literature repository access restrictions and computational resource costs Wang et al. (2022b); Chen et al. (2024). The current instruction fine-tuning data construction method Wang et al. (2023); Zhang et al. (2023), which relies heavily on large model generation, amplifies discrepancies between response styles and human preferences, posing a significant challenge. This paper introduces the first energy domain-specific large language model, a novel development that completes full-process training, constructs the domain’s first training and evaluation datasets, and enhances model response alignment with human preferences through RLHF.
@@ 3.1 Pre-training Data: EnerBase
In this work, we conducted specialized text data pre-training based on the Qwen2.5-7B foundation model. The constructed Smart Energy Corpus includes open-access academic preprints, authoritative journal papers, specialized publications, domain-specific modeling toolkits, and application code, and IEA energy datasets from the smart energy domain. Data was collected from arXiv, Web of Science (WoS), Acemap, Github, and HuggingFace platforms. After data preprocessing, we constructed EnerBase, a high-quality pre-training corpus of 3 billion tokens to enhance the model’s accumulation of professional knowledge and technical application capabilities in the smart energy domain. In brief, the statistical characteristics of Smart Energy Corpus are shown in Table 1.
3.1.1 Smart Energy Corpus Collection.
Scientific Literature. The smart energy domain’s extensive scientific literature provides a high-quality training corpus for large language models, enhancing their domain-specific knowledge understanding and application capabilities. To ensure the comprehensiveness of the corpus, we systematically decomposed the smart energy domain into 14 specialized sub-domains, including load forecasting and energy storage, and collected data for each separately. 1) Open-access Academic Preprints (OAP): We crawled 173,541 PDF files from arXiv using subdomain-specific keywords, establishing the quantitative foundation of our Smart Energy Corpus. 2) Open-access Authoritative Journal Papers (OAJP): We extracted metadata from WoS for leading energy journals and crawled 32,459 PDF files, establishing the qualitative foundation of our Smart Energy Corpus. 3) Specialized Publications (SP): We crawled 363 professional book PDF files from the ACEmap platform, enriching the Smart Energy Corpus knowledge framework.
@@
International Energy Agency Datasets (IEAD). The IEA, covering 75% of global energy demand, has evolved from an oil crisis response mechanism to a platform governing energy security, economic growth, and environmental protection. Its statistics system provides authoritative data on supply-demand balance, emissions, renewables, and efficiency indicators across 170+ countries. To enhance LLMs’ analytical capabilities for energy transition assessment, we incorporated the IEA_Energy_Dataset Li (2023) with 358,446 data points into our training corpus.
@@ 3.1.2 Smart Energy Corpus Processing
PDF Convert. Our Smart Energy Corpus primarily exists in PDF format, necessitating conversion to a unified format suitable for model training. Scientific literature contains abundant structured information, including tables, equations, and formulas; direct conversion to TXT format would result in critical information loss, causing large language models to learn incomplete or incorrect content. Therefore, we selected Markdown as our unified conversion format to preserve these essential structural elements.
To balance computational throughput with structural integrity, we developed Python scripts based on Marker Paruchuri (2025). For processing efficiency, we deployed 10 servers equipped with NVIDIA RTX 4090 GPUs in a distributed architecture, each server configured with six parallel conversion workers. To enhance quality, we integrated OpenAI’s GPT-4o as an intelligent agent (Parsing Agent) to perform table reconstruction, mathematical formula standardization, form parsing, figure description generation, and reference normalization, ensuring structural completeness. Detailed hyperparameter configurations are provided in the supplementary table. Our system achieved an average processing speed of 2.21 seconds per page, completing the entire conversion process within 5 days. Fig. 2 demonstrates sample conversion results. The computationally efficient and structurally complete PDF-to-Markdown conversion framework, based on intelligent agents, presented in this paper, has been open-sourced on GitHub along with the dataset.
@@
Deduplication. Nevertheless, the Smart Energy Corpus inevitably contains a proportion of semantically similar fragments, causing the model during pre-training to update along nearly identical gradient directions and thus to “memorise” specific passages rather than acquire generalisable logical patterns Tirumala et al. (2023). To address this problem, following the methodology outlined in Abbas et al. (2023), we developed an efficient large-scale deduplication agent, Corpus Distiller, built on BERT-base. Corpus Distiller first performs K-Means clustering in the embedding space and subsequently removes samples located within the same epsilon-ball in each cluster.
@@ 3.2 Instruction Tuning Data
Instruction Tuning is the key to bridging large‑scale unsupervised pre‑trained models with downstream applications. We have constructed a two‑phase instruction fine‑tuning framework of “Universal Human Instruction Comprehension (UHIC) to Domain-specific Task Adaptation (DS-TA)”: first, high‑quality general instruction samples are employed to conduct preliminary fine‑tuning, enabling the model to learn to accomplish tasks according to natural‑language instructions; subsequently, knowledge‑intensive, specialized data are introduced for further fine‑tuning, thereby enhancing the model’s adaptability to domain‑specific tasks. For these two phases, we curate a complementary general instruction dataset and knowledge‑intensive dataset EnerInstruct, each uniformly organized in an <instruction,input,output> triplet format.

### v2

@@ Abstract
In the pursuit of carbon neutrality, smart energy systems integrating renewable energy, storage, and demand response have become central to energy system transformation. However, fragmented and rapidly evolving interdisciplinary knowledge increases the cognitive and information integration burden for operational decision-making. Although large language models (LLMs) have shown promise in smart energy tasks through fine-tuning or prompt engineering, their lack of domain knowledge and physical constraints often results in semantically plausible but physically inconsistent outputs, limiting their engineering reliability. To address these challenges, we introduce Helios, the first large language model tailored to the smart energy domain, together with a comprehensive suite of resources to advance LLM research in this field. Specifically, we develop EnerSys, a multi-agent collaborative framework for end-to-end dataset construction, through which we produce: (1) the first smart energy knowledge base, EnerBase, to enrich the model’s foundational expertise; (2) the first instruction tuning dataset, EnerInstruct, to strengthen performance on domain-specific downstream tasks; and (3) the first Reinforcement Learning from Human Feedback (RLHF) dataset, EnerReinforce, to align the model with human preferences and industry standards. Leveraging these resources, Helios undergoes large-scale pretraining, instruction tuning, and RLHF. We also release EnerBench, the first benchmark for evaluating LLMs in smart energy scenarios, and demonstrate that our approach significantly enhances domain knowledge mastery, task execution accuracy, and alignment with human preferences.
@@
Haoyu Jiang1,†thanks: These authors contributed equally to this work.  Fanjie Zeng2,†   Boan Qu2,†   Xiaojie Lin1,∗thanks: Corresponding author.  Wei Zhong1,3 1 College of Energy Engineering, Zhejiang University 2 Polytechnic Institute, Zhejiang University 3Shanghai Institute for Advanced Study, Zhejiang University {haoyu.jiang, zengfanjie, boan.qu, xiaojie.lin, wzhong}@zju.edu.cn https://helios-llm.github.io/
@@ 1 Introduction
Figure 1: Limitations of Scheduling Decision-Making in the Smart Energy Domain.
@@
Driven by the global pursuit of carbon neutrality, smart energy systems must enhance overall efficiency through the intelligent coordination of renewable energy integration, energy storage dispatch, and demand‑side response Lund et al. (2017). Smart energy is highly interdisciplinary, encompassing power engineering, information science, economics, and other fields, and its knowledge base is fragmented and rapidly evolving Ceglia et al. (2020); Thellufsen et al. (2020). Building on recent advances in general large language models (LLMs) in semantic understanding, logical reasoning, and multitask generalization, a growing body of research has used fine‑tuning and prompt engineering to adapt LLMs to task‑specific applications in smart energy, such as load forecasting Jin et al. (2023); Liao et al. (2025); Hu et al. (2025), building energy consumption modeling Wang et al. (2025b); Jiang et al. (2024), and HVAC fault diagnosis Zhang et al. (2025), thereby supporting case modeling and intelligent decision making.
@@
However, general LLMs often deliver reasoning that is semantically plausible yet physically invalid Friel and Sanyal (2023). This limitation arises chiefly because their pre‑training corpora lack reliable knowledge from the smart energy domain, leaving the models without essential domain context and physical constraints Deng et al. (2024). Current approaches Hu et al. (2025); Zhang et al. (2025) mainly invoke the prior knowledge already embedded in LLMs and do not explicitly enrich them with smart energy expertise. To alleviate these challenges, we introduce the first-ever open‑sourced foundational LLM for the smart‑energy domain, referred to as Helios (Originating from the ancient Greek sun‑god, signifies the illumination of the pathway toward sustainable development through the radiance of smart energy, thereby advancing the harmonious co‑existence of humanity and the natural environment). Helios is capable of effectively tackling a broad spectrum of smart‑energy tasks. Furthermore, we present EnerSys, an end-to-end multiagent collaborative framework for dataset construction that integrates automated data generation, screening, and refinement, thereby furnishing Helios with an extensive and high‑quality data foundation.
@@
EnerSys covers three dataset‑construction phases (as shown in Fig. 1): In the construction of the pre-training dataset, the Parsing‑Agent and Deduplication Agent extract structured knowledge from the Smart Energy Corpus (scientific papers, domain‑modeling code, IEA datasets, etc.) and eliminate redundancy, building a comprehensive, balanced smart energy domain knowledge base, EnerBase; In instruction‑tuning dataset construction, on expert‑crafted seed data, we deploy Expert‑Agents for each of 14 smart‑energy sub‑domains, letting them generate instruction–response pairs from the seeds and a high‑quality corpus; the Check‑Agent then scores samples on accuracy, completeness, relevance, and usability, and the Refine‑Agent automatically fixes those below par. This pipeline yielded the EnerInstruct; In the RLHF dataset construction, agents like Write‑like‑Human craft multi‑level candidate answers to given questions, thereby creating the EnerReinforce to supply the reward model with differentiated contrastive samples. Using these datasets, we complete Helios pre‑training (adding domain basics), supervised fine‑tuning (boosting downstream skills), and RLHF reinforcement (aligning with human preferences). Concurrently, adhering to a dual‑track paradigm of “public item‑bank retrieval + expert‑targeted design,” we build EnerBench, containing 625 subjective and 887 objective questions, to systematically assess LLMs performance in smart‑energy scenarios. Experiments show Helios surpasses general‑purpose LLMs on both tasks, with output style tightly matching professional context. Our contributions can be summarized as follows:
@@ 2 Related Work
Foundation Language Models. LLMs trained on vast amounts of diverse and heterogeneous data, have accumulated extensive domain knowledge and contextual modeling capabilities. They have demonstrated human-level performance in many tasks. LLMs can be categorized into two types: 1) Closed-source models (such as OpenAI o1 Jaech et al. (2024) and Claude): These models provide inference interfaces via APIs, making them suitable for industrial-grade deployment without the need for building custom computational resources. However, they cannot be customized or extended according to specific needs; 2) Open-source models (such as DeepSeek Guo et al. (2025); DeepSeek-AI et al. (2024), Llama and Qwen Bai et al. (2023, 2025)): These models offer complete training weights, allowing for customized applications based on downstream task requirements. This has led to the development of instruction-tuned models like Alpaca Taori et al. (2023a), Vicuna Chiang et al. (2023), and Dolly Conover et al. (2023a). In this process, the quality of datasets becomes a critical factor affecting training outcomes.
@@
Domain Language Models. LLMs excel in general reasoning Jiang et al. (2025a); Nam et al. (2024), their performance in specialized applications is hampered by a lack of domain expertise. This limitation has led researchers to adapt foundation models for vertical domains such as medicine Tian et al. (2024); Lin et al. (2025a), chemistry Zhang et al. (2024); Zheng et al. (2025), ocean science Bi et al. (2024), and geography Deng et al. (2024). However, most domains are still in the early stages of exploration. In the energy sector, existing research primarily leverages general models’ prior knowledge through prompt engineering or fine-tuning for load forecasting Jin et al. (2023); Jiang et al. (2025b); Wu and Ling (2024); Wang et al. (2025a), building energy modeling Wang et al. (2025b); Jiang et al. (2024), teaching assistant Lin et al. (2025b), and HVAC fault diagnosis Zhang et al. (2025). These approaches focus on application (applying LLMs’ prior knowledge to downstream tasks) rather than accumulation (enriching models with energy domain knowledge through pretraining). Furthermore, the energy domain faces a scarcity of high-quality training data due to literature repository access restrictions and computational resource costs Wang et al. (2022b); Chen et al. (2024). The current instruction fine-tuning data construction method Wang et al. (2023); Zhang et al. (2023), which relies heavily on large model generation, amplifies discrepancies between response styles and human preferences, posing a significant challenge. This paper introduces the first energy domain-specific large language model, a novel development that completes full-process training, constructs the domain’s first training and evaluation datasets, and enhances model response alignment with human preferences through RLHF.
Multi-Agent Systems. Due to the extensive domain knowledge and robust semantic understanding capabilities of LLMs, they are employed as core components of agents to support intelligent decision-making and natural language interaction Guo et al. (2024). In single-agent systems, a single agent carries out decision-making and task execution, which is suitable for structured scenarios with fewer variables Chen et al. (2023). However, as the complexity of problems increases, single-agent systems face issues such as low decision-making efficiency, slow response times, and poor fault tolerance Amirkhani and Barshooi (2022). In contrast, multi-agent systems can effectively address these challenges through the collaboration of specialized agents and have been widely applied in complex scenarios such as interactive games Mao et al. (2023); Xu et al. (2023), financial markets Li et al. (2023), and social simulations Park et al. (2023). Currently, some scholars Bi et al. (2024); Ni and Buehler (2024) are exploring ways to improve the efficiency and representativeness of domain dataset construction through multi-agent collaboration and distributed decision-making mechanisms. However, constructing domain datasets typically involves multiple steps, including data generation, deduplication, filtering, and optimization. Existing research often focuses only on optimizing specific steps and has not fully leveraged the potential of multi-agent systems in dataset construction.
@@ 3.1 Pre-training Data: EnerBase
In this work, we conducted specialized text data pre-training based on the Qwen2.5-7B foundation model. The constructed Smart Energy Corpus includes open-access academic preprints, authoritative journal papers, specialized publications, domain-specific modeling toolkits, and application code, and IEA energy datasets from the smart energy domain. Data was collected from arXiv, Web of Science (WoS), Acemap, Github, and HuggingFace platforms. After data preprocessing, we constructed EnerBase, a high-quality pre-training corpus of approximately 3 billion tokens to enhance the model’s accumulation of professional knowledge and technical application capabilities in the smart energy domain. In brief, the statistical characteristics of Smart Energy Corpus are shown in Table 1.
3.1.1 Smart Energy Corpus Collection
Scientific Literature. The smart energy domain’s extensive scientific literature provides a high-quality training corpus for LLMs, enhancing their domain-specific knowledge understanding and application capabilities. To ensure the comprehensiveness of the corpus, we systematically decomposed the smart energy domain into 14 specialized sub-domains, and collected data for each separately. 1) Open-access Academic Preprints (OAP): We crawled 173,541 PDF files from arXiv using subdomain-specific keywords, establishing the quantitative foundation of our Smart Energy Corpus. 2) Open-access Authoritative Journal Papers (OAJP): We extracted metadata from WoS for leading energy journals and crawled 32,459 PDF files, establishing the qualitative foundation of our Smart Energy Corpus. 3) Specialized Publications (SP): We crawled 363 book PDF files from the Acemap, enriching the Smart Energy Corpus knowledge framework.
@@ 3.1.1 Smart Energy Corpus Collection
International Energy Agency Datasets (IEAD). The IEA, covering 75% of global energy demand, has evolved from an oil crisis response mechanism to a platform governing energy security, economic growth, and environmental protection. Its statistics system provides authoritative data on supply-demand balance, emissions, renewables, and efficiency indicators across 170+ countries. To enhance LLMs’ analytical capabilities for energy transition assessment, we incorporated the IEA_Energy_Dataset Li (2023) with 358,466 data points into our training corpus.
@@ 3.1.2 Smart Energy Corpus Processing
PDF Convert. Collected corpus primarily exists in PDF format, necessitating conversion to a unified format suitable for model training. Scientific literature contains abundant structured information, including tables, equations, and formulas; direct conversion to TXT format would result in critical information loss, causing LLMs to learn incomplete or incorrect content. Therefore, we selected Markdown as our unified conversion format to preserve these essential structural elements.
To balance computational throughput with structural integrity, we developed Python scripts based on Marker Paruchuri (2025). For processing efficiency, we deployed 10 servers equipped with NVIDIA RTX 4090 GPUs in a distributed architecture, each server configured with six parallel conversion workers. To enhance quality, we integrated OpenAI’s GPT-4o as an intelligent agent (Parsing-Agent) to perform table reconstruction, mathematical formula standardization, form parsing, figure description generation, and reference normalization, ensuring structural completeness. Detailed hyperparameter configurations are provided in the supplementary table. Our system achieved an average processing speed of 2.21 seconds per page, completing the entire conversion process within 5 days. Fig. 3 demonstrates sample conversion results. The computationally efficient and structurally complete PDF-to-Markdown conversion framework, based on intelligent agents, presented in this paper, has been open-sourced on GitHub along with the dataset.
@@
Deduplication. Nevertheless, the Smart Energy Corpus inevitably contains a proportion of semantically similar fragments, causing the model during pre-training to update along nearly identical gradient directions and thus to "memorise" specific passages rather than acquire generalisable logical patterns Tirumala et al. (2023). To address this problem, following the methodology outlined in Abbas et al. (2023), we developed an efficient large-scale deduplication agent, Corpus Distiller, built on BERT-base. Corpus Distiller first performs K-Means clustering in the embedding space and subsequently removes samples located within the same epsilon-ball in each cluster.
Table 2: Datasets used to train Helios during the Universal Human Instruction Comprehension phase.
Dataset Prompts
Alpaca-cleaned Taori et al. (2023b) 51 800
Dolly-15K Conover et al. (2023b) 15 011
Natural-Instructions Muennighoff (2022) 30 000
python_code_25k FLOCK4H (2023) 24 813
OpenR1-Math-220k R1 (2025) 28 120
Toolbench Qin et al. (2023) 10 328
Total 160 072
@@ 3.2 Instruction Tuning Data
Instruction Tuning is the key to bridging large‑scale unsupervised pre‑trained models with downstream applications. We have constructed a two‑phase instruction fine‑tuning framework of "Universal Human Instruction Comprehension (UHIC) to Domain-specific Task Adaptation (DS-TA)": First, high‑quality general instruction samples are employed to conduct preliminary fine‑tuning, enabling the model to learn to accomplish tasks according to natural‑language instructions; subsequently, knowledge‑intensive, specialized data are introduced for further fine‑tuning, thereby enhancing the model’s adaptability to domain‑specific tasks. For these two phases, we curate a complementary general instruction dataset and knowledge‑intensive dataset EnerInstruct, each uniformly organized in an <instruction,input,output> triplet format.